OpenStudy (anonymous):

Two angles form a linear pair. The measure of one angle is four times the measure of the other angle. Find the measure of each angle. Please help! I have no idea what to do! >.<

OpenStudy (anonymous):

Ok, so we have two angles: \(a\) and \(b\). They form a linear pair so: \(a+b = 180^\circ \) We also know that \(a = 2b\) since one is twice the size of the other. Thus we substitute \(2b\) into \(a\) in the first equation to get: \[2b + b = 180^\circ \] \[3b = 180^\circ \] \[b = 60^\circ \] Then we substitue \(60^\circ \) into \(b\) in our second equation: \[a+60^\circ =180^\circ \] \[a =120^\circ \] So there we have it. \(a=120^\circ , b=60^\circ\)

OpenStudy (anonymous):

Actually, this is the case where one is twice the size of the other. So it should be \(a=4b\) for our second equation. Sorry. But think you can do it?

OpenStudy (anonymous):

I can redo it the correct way if you'd like. But you should try if yourself first.

OpenStudy (anonymous):

a+b= 180 a=4b 3b+b=180 4b=180 b=45 ?

OpenStudy (anonymous):

then substitute?

OpenStudy (anonymous):

No a+b= 180 a=4b 4b+b=180 <--- this is what it should be here. Good work though, you're getting there.

OpenStudy (anonymous):

b=36, a=144 ?